# SAM-Assisted Remote Sensing Imagery Semantic Segmentation with Object and Boundary Constraints

### 05 Dec 2023 v1

@@ Abstract
Semantic segmentation of remote sensing imagery plays a pivotal role in extracting precise information for diverse down-stream applications. Recent development of the Segment Anything Model (SAM), an advanced general-purpose segmentation model, has revolutionized this field, presenting new avenues for accurate and efficient segmentation. However, SAM is limited to generating segmentation results without class information. Consequently, the utilization of such a powerful general vision model for semantic segmentation in remote sensing images has become a focal point of research. In this paper, we present a streamlined framework aimed at leveraging the raw output of SAM by exploiting two novel concepts called SAM-Generated Object (SGO) and SAM-Generated Boundary (SGB). More specifically, we propose a novel object loss and further introduce a boundary loss as augmentative components to aid in model optimization in a general semantic segmentation framework. Taking into account the content characteristics of SGO, we introduce the concept of object consistency to leverage segmented regions lacking semantic information. By imposing constraints on the consistency of predicted values within objects, the object loss aims to enhance semantic segmentation performance. Furthermore, the boundary loss capitalizes on the distinctive features of SGB by directing the model’s attention to the boundary information of the object. Experimental results on two well-known datasets, namely ISPRS Vaihingen and LoveDA Urban, demonstrate the effectiveness of our proposed method. The source code for this work will be accessible at https://github.com/sstary/SSRS.
@@
Remote Sensing, Semantic Segmentation, SAM, Object Loss, Boundary Loss
@@ I Introduction
Semantic segmentation of remote sensing imagery entails assigning semantic labels to individual pixels within images captured by various remote sensing sensors. This process is fundamental for diverse downstream geoscience applications, including environmental monitoring [1], land cover mapping [2], and disaster management [3, 4]. The primary objective is to accurately partition the image into distinct regions representing different semantic classes, facilitating automated analysis and interpretation of remote sensing data. The advent of deep learning techniques [5, 6, 7, 8] has ushered in numerous high-performance, problem-specific methods in this field. According to the type of network, these methods can be primarily categorized into convolutional neural network (CNN) [9, 10, 11], transformer-based approaches [12, 13] and hybrid architecture [14, 15, 16, 17] for remote sensing applications.
@@
Recently, a notable foundation model called Segment Anything Model (SAM) [18] designed for image segmentation gains considerable attention from the computer vision community. Trained on an extensive dataset of $11$ million natural images and over one billion masks, SAM distinguishes itself by enabling zero-shot segmentation of new visual objects without prior exposure to them. As the first foundation model for general image segmentation, SAM and its original paper have garnered over $1000$ citations within six months according to the google scholar website. However, two conspicuous limitations hinder the application of SAM to remote sensing image semantic segmentation tasks. Firstly, the generated segmentation masks lack semantic labels. Secondly, due to the disparities between natural images and remote sensing images, SAM’s effectiveness in remote sensing tasks is compromised [19].
@@
To address these limitations and enhance SAM’s performance in remote sensing semantic segmentation tasks, researchers have explored various methods from different perspectives. These approaches include adapting or enhancing SAM [20, 21, 22], employing few-shot or zero-shot learning strategies [23, 24, 25, 26] and leveraging prompt learning techniques [27, 28, 29, 30, 31]. In particular, SAM-CD [21] directly utilizes FastSAM [32] to extract image features while incorporating the underlying temporal constraints in remote sensing images for change detection supervision. RS-CLIP [24] employs a curriculum learning strategy to boost the performance of zero-shot classification of remote sensing images using SAM through multiple stages of model finetuning. SAMRS [28] efficiently generates a large-scale remote sensing segmentation dataset by employing various prompts, thereby significantly expanding existing remote sensing datasets. However, these methods often necessitate the artificial design of complex fine-tuning mechanisms or prompt learning strategies. The former involves modifying the structures of the general semantic segmentation model and training strategies, while the latter requires specific prompts tailored to different datasets. Unfortunately, these factors impede the seamless integration of SAM into the field of remote sensing.
@@
Figure 1: Visual examples. (a) Images, (b) Ground truth, (c) SGO, (d) SGB. The first two rows show samples of size $512\times 512$ from ISPRS Vaihingen. The last two rows show samples of size $1024\times 1024$ from LoveDA Urban. It can be observed that SGO and SGB provide a wealth of detailed information about ground objects from both object and boundary perspectives.
@@
To cope with the aforementioned challenges, we develop a simple yet effective framework to utilize SAM from the perspectives of both objects and boundaries. Our observations reveal that SAM encounters challenges in accurately segmenting remote sensing images due to disparities between remote sensing images and natural images. Nevertheless, SAM exhibits proficiency in recognizing objects, as evident in Fig. 1. It shows that SAM-Generated Object (SGO) and SAM-Generated Boundary (SGB) can provide detailed object and boundary information. To fully leverage their potential while minimizing modifications to the general semantic segmentation model, we propose a novel loss function, namely object loss and further introduce a boundary loss to aid in model training.
@@
For the object loss, it is observed that the regions in SGO as shown in Fig. 1(c) are actually objects without semantic information. This insight motivates the design of the object loss, aiming to ensure consistency within a segmented object. Furthermore, we introduce the boundary loss [33] to encourage the segmentation model to better consider segment boundaries based on the detailed SGB information. As both loss functions take the semantic segmentation output from the general segmentation model as inputs, there is no need for additional segmentation heads at the tail of the decoder. By harnessing a simple and direct utilization of SGO and SGB, the foundational visual knowledge embedded in SAM contributes to enhancing the semantic segmentation performance of remote sensing images. This approach holds the potential for direct implementation across various tasks involving the combination of SAM and remote sensing. The contributions of this work can be summarized as follows:
@@
  • We have developed a streamlined framework that efficiently leverages two novel concepts called SGO and SGB for remote sensing segmentation segmentation, which underscores the value and effectiveness of the raw output of SAM. Notably, our approach distinguishes itself by not necessitating specific designs for the model, training strategy, or pseudo-label generation, in contrast to other existing SAM-based methods employed in remote sensing;
@@
  • We propose a novel object loss and further introduce the boundary loss to assist with model optimization, making full use of the raw output of SAM. This is achieved without relying on semantic information, focusing on two crucial perspectives: object and boundary. To our best knowledge, the object loss is the first method in semantic segmentation tasks that can directly utilize the raw output of SAM without the need for additional class prompts;
@@
  • Extensive experiments on two well-known, publicly available remote sensing datasets, ISPRS Vaihingen and LoveDA Urban, confirm that the proposed approach can be widely adapted to semantic segmentation tasks of different datasets and different models. We believe that this approach has the potential to significantly expand the application of SGO and SGB, unlocking the full capabilities of large models like SAM in remote sensing semantic segmentation tasks.
@@
The remainder of this paper is organized as follows. Sec. II first reviews the related works of SAM in dirrerent fields. After that, Sec. III presents the proposed method in detail, whereas Sec. IV provides experimental results and discussions. Finally, the conclusion is given in Sec. V.
Figure 2: The schematic diagram of the proposed method. The black arrows describe general semantic segmentation methods. The image $X$ is fed into the segmentation model that generates the segmentation output $P$ before the segmentation loss $L_{seg}$ is computed by segmentation output and ground truth $Y$, which is used for back-propagation to update the model. The red arrows represent the processes we designed to leverage SGO and SGB. Two additional loss functions $L_{obj}$ and $L_{bdy}$ are computed by SGO $Y_{o}$ and SGB $Y_{b}$ that will also be used for model optimization together with $L_{seg}$. It should be noted that $P\in\mathbb{R}^{H\times W\times C}$ in this figure is the predicted score of the model for $C$ classes. Here we take the category label corresponding to the maximum score of each pixel for visualization.
@@ II-A Segment Anything Model
The Segment Anything Model (SAM), introduced by Meta AI [34], is a large vision transformer [8]-based model trained on a substantial visual corpus. It is a foundation model aiming to address specific downstream image segmentation tasks. One of the challenges in applying deep neural networks to real-world semantic segmentation applications, is the need for large amounts of well-annotated training data. SAM is an effective tool to address this issue by enabling zero-shot generalization to unseen images and objects based on user-provided prompts. The framework of SAM has three main components, namely an image encoder, a prompt encoder, and a mask decoder. The image encoder utilizes a vision transformer-based approach for extracting image features. The prompt encoder incorporates user interactions for segmentation tasks. Different types of prompts, including points, text, box, and mask prompts, are supported. The mask decoder comprises transformer layers with dynamic mask prediction heads and an Intersection-over-Union (IoU) score regression head. It maps the encoder embedding, prompt embeddings, and an output token to a mask. For user convenience, SAM offers a set of APIs with which segmentation masks can be obtained in just a few lines of code. Different segmentation mode options, i.e., fully automatic, bounding box, and point mode, are supported for different prompts in APIs.
Currently, SAM has made significant strides in diverse fields. In the realm of medical image processing, [35] introduced a straightforward image enhancement method by combining SAM-generated masks with raw images. Additionally, nnSAM [36] integrated the encoder of UNet with the pre-trained SAM encoder, harnessing the feature extraction capabilities of this large foundational vision model. Furthermore, [37, 38] delved into exploring SAM’s capacity to generate pseudo labels. Through these distinct technical approaches, these methods have propelled the application and development of SAM across various domains.
@@ III Methodology
The schematic diagram of the proposed framework is depicted in Fig. 2. In the conventional semantic segmentation task, the image is fed into the segmentation model to generate segmentation output, which is then used to calculate the segmentation loss. Subsequently, the model is updated based on the back-propagation algorithm, as depicted by the black arrows in Fig. 2. The additional steps introduced in our framework are highlighted by red arrows. Initially, we directly generate SGO and SGB using the SAM provided by Meta AI [18]. These outputs play a crucial role in computing the object loss and boundary loss, respectively, thereby contributing to the model’s training. In this section, we will provide detailed explanations on SAM-Generation, network training, and the associated loss functions.
III-A SAM-Generation
SAM provides a grid prompt technique to automatically process images. Given the input remote sensing image denote by $X\in\mathbb{R}^{H\times W\times 3}$, SAM can generate segmentation masks across the entire image at all plausible locations in the grid prompt setting [35]. In this study, we refer to the segmentation mask as the object, considering each segmentation mask as an individual enclosed region that can be regarded as an object. The generated objects are then stored into a list where we set a threshold of $K$ to limit the maximum number of objects in $X$. Meanwhile, we also establish a threshold $S$ to set a minimum limit on the number of pixels that a single object can contain, effectively filtering out very small segmentation masks. Consequently, a SAM-Generation object, denoted as SGO and represented by $Y_{o}\in\mathbb{R}^{H\times W}$ can be obtained, where the value of each pixel falls within the range of $[0,K]$. Pixels not segmented as objects, as well as boundaries, are assigned a value of 00, while the objects in $Y_{o}$ are indexed by an identifier denoted as $i$, where $i\in[1,K]$. Concurrently, a boundary prior map is derived from SGO. This process involves outlining the exterior boundaries of each object within the list and merging these boundaries to produce a comprehensive boundary prior map, namely SGB, denoted by $Y_{b}\in\mathbb{R}^{H\times W}$. Unless specified otherwise, the boundaries in $Y_{b}$ are set to $255$, while others are set to 00 in the sequel. Visual examples of SGO and SGB are illustrated in Fig. 1 (c) and (d).
@@ III-B Network Training
The classical encoder-decoder networks, e.g., UNetformer [14], are widely used in semantic segmentation methods. In this work, we use them as the segmentation model in the proposed framework shown in Fig. 2. Given the input image $X$, the segmentation model produces the predicted segmentation output denote by $P\in\mathbb{R}^{H\times W\times C}$ where $C$ is the number of categories of ground objects. The learning objective is to minimize the following cross entropy-based segmentation loss with respect to the parameters of segmentation model:
@@
where $L_{obj}$ and $L_{bdy}$ stand for the object loss and boundary loss, respectively, while $\lambda_{o}$ and $\lambda_{b}$ are two weighting coefficients to balance them.
III-B1 Object Loss
The object loss is designed to maintain pixel consistency within the objects in a given input image. Given the input $X$, the output of the segmentation model is denoted as $P$. To compute the object loss, we iterate through all the objects in $Y_{o}$ in order of id $i$. For each object, we first extract its mask $M^{i}$, i.e. the region in $Y_{o}$ where the pixel value equals $i$. The object feature is then obtained by:

### 20 Dec 2023 v2

@@ Abstract
Semantic segmentation of remote sensing imagery plays a pivotal role in extracting precise information for diverse downstream applications. Recent development of the Segment Anything Model (SAM), an advanced general-purpose segmentation model, has revolutionized this field, presenting new avenues for accurate and efficient segmentation. However, SAM is limited to generating segmentation results without class information. Meanwhile, the segmentation map predicted by current methods generally exhibit excessive fragmentation and inaccuracy of boundary. This paper introduces a streamlined framework designed to leverage the raw output of SAM by exploiting two novel concepts called SAM-Generated Object (SGO) and SAM-Generated Boundary (SGB). More specifically, we propose a novel object consistency loss and further introduce a boundary preservation loss in this work. Considering the content characteristics of SGO, we introduce the concept of object consistency to leverage segmented regions lacking semantic information. By imposing constraints on the consistency of predicted values within objects, the object consistency loss aims to enhance semantic segmentation performance. Furthermore, the boundary preservation loss capitalizes on the distinctive features of SGB by directing the model’s attention to the boundary information of the object. Experimental results on two well-known datasets, ISPRS Vaihingen and LoveDA Urban, demonstrate the effectiveness and broad applicability of our proposed method. The source code for this work will be accessible at https://github.com/sstary/SSRS.
@@
Remote Sensing, Semantic Segmentation, SAM, Object Consistency Loss, Boundary Preservation Loss
@@ I Introduction
Semantic segmentation of remote sensing imagery entails assigning semantic labels to individual pixels within images captured by various remote sensing sensors. This process is fundamental for diverse downstream geoscience applications, including environmental monitoring [1, 2], land cover mapping [3, 4], and disaster management [5, 6, 7, 8]. The primary objective is to accurately partition the image into distinct regions representing different semantic classes, facilitating automated analysis and interpretation of remote sensing data. The advent of deep learning techniques [9, 10, 11, 12] has ushered in numerous high-performance, problem-specific methods in this field. According to the type of network, these methods can be primarily categorized into convolutional neural network (CNN) [13, 14, 15, 16], transformer-based approaches [17, 18, 19, 20] and hybrid architecture [21, 22, 23, 24] for remote sensing applications.
@@
Recently, a notable foundation model called Segment Anything Model (SAM) [25] designed for image segmentation has gained considerable attention from the computer vision community. Trained on an extensive dataset of $11$ million natural images and over one billion masks, SAM distinguishes itself by enabling zero-shot segmentation of new visual objects without prior exposure to them. As the first foundation model for general image segmentation, SAM and its original paper have garnered over $1000$ citations within six months, according to the Google Scholar website. However, two conspicuous limitations hinder the application of SAM to remote sensing image semantic segmentation tasks. Firstly, the generated segmentation masks lack semantic labels. Secondly, due to the disparities between natural images and remote sensing images, SAM’s effectiveness in remote sensing tasks is compromised [26].
@@
Researchers have explored various methods from different perspectives to address these limitations and enhance SAM’s performance in remote sensing image semantic segmentation tasks. These approaches include adapting or enhancing SAM [27, 28, 29], employing few-shot or zero-shot learning strategies [30, 31, 32, 33] and leveraging prompt learning techniques [34, 35, 36, 37, 38, 39, 40]. In particular, SAM-CD [28] directly utilizes FastSAM [41] to extract image features while incorporating the underlying temporal constraints in remote sensing images for change detection supervision. RS-CLIP [31] employs a curriculum learning strategy to boost the performance of zero-shot classification of remote sensing images using SAM through multiple stages of model finetuning. SAMRS [35] efficiently generates a large-scale remote sensing segmentation dataset by employing various prompts, thereby significantly expanding existing remote sensing datasets. However, these methods often necessitate the artificial design of complex fine-tuning mechanisms or prompt learning strategies. The former involves modifying the structures of the general semantic segmentation model and training strategies, while the latter requires specific prompts tailored to different datasets. Unfortunately, these factors impede the seamless integration of SAM into the field of remote sensing.
@@
Figure 1: Visual examples. (a) Images, (b) Ground truth, (c) SGO, (d) SGB. The first two rows show samples of size $512\times 512$ from ISPRS Vaihingen. The last two rows show samples of size $512\times 512$ from LoveDA Urban. It can be observed that SGO and SGB provide a wealth of detailed information about ground objects from both object and boundary perspectives.
@@
To cope with the aforementioned challenges, we develop a simple yet effective framework to utilize SAM from the perspectives of both objects and boundaries. Our observations reveal that SAM encounters challenges in accurately segmenting remote sensing images due to disparities between remote sensing images and natural images. Nevertheless, SAM exhibits proficiency in recognizing objects, as evident in Fig. 1. It shows that SAM-Generated Object (SGO) and SAM-Generated Boundary (SGB) can provide detailed object and boundary information. To fully leverage their potential while minimizing modifications to the general semantic segmentation model, we propose a novel loss function, namely object consistency loss and further introduce a boundary preservation loss to aid in model training.
@@
For the object consistency loss, it is observed that the regions in SGO, as shown in Fig. 1 (c), are actually objects without semantic information. This insight motivates the design of the object consistency loss, aiming to ensure consistency within a segmented object. Furthermore, we introduce the boundary preservation loss [42] to encourage the semantic segmentation model to consider better segment boundaries based on the detailed SGB information. As both loss functions take the semantic segmentation output from the general semantic segmentation model as inputs, there is no need for additional segmentation heads at the tail of the decoder. By harnessing a simple and direct utilization of SGO and SGB, the foundational visual knowledge embedded in SAM contributes to enhancing the semantic segmentation performance of remote sensing images. This approach holds the potential for direct implementation across various tasks and semantic segmentation models involving the combination of SAM and remote sensing. The contributions of this work can be summarized as follows:
@@
  • A streamlined framework is proposed to efficiently leverage two novel concepts called SGO and SGB for remote sensing image semantic segmentation, which emphasizes the value and effectiveness of the raw output of SAM. Notably, our approach stands out by not necessitating specific designs for the semantic segmentation model, training strategy, or pseudo-label generation, in contrast to other existing SAM-based methods employed in remote sensing;
@@
  • We propose a novel object consistency loss by restraining the consistency of pixels within an object, and further introduce the boundary preservation loss to assist with model optimization, making full use of the raw output of SAM. This is achieved without relying on semantic information, focusing on two crucial perspectives: object and boundary. To our best knowledge, this work is the first to introduce object and boundary constraints into semantic segmentation tasks to refine the segmentation results by directly utilizing the raw output of SAM without the need for additional class prompts;
@@
  • Extensive experiments on two well-known, publicly available remote sensing datasets, ISPRS Vaihingen and LoveDA Urban, and four representative semantic segmentation models confirm that the proposed approach can be widely adapted to semantic segmentation tasks of different datasets and different general models. We believe that this approach has the potential to significantly expand the application of SGO and SGB, unlocking the full capabilities of large models like SAM in remote sensing image semantic segmentation tasks.
@@
The remainder of this paper is organized as follows. Section II first reviews the related works of SAM in different fields. After that, Section III presents the proposed method in detail, whereas Section IV provides experimental results and discussions. Finally, the conclusion is given in Section V.
Figure 2: Comparison of frameworks between SAM and our approach. (a) Schematic of the SAM framework [25]. (b) Schematic of the proposed SAM-based pre-processing approach. All parameters of SAM are frozen, and SGO and SGB are generated with only points prompt.
@@ II-A Segment Anything Model
The Segment Anything Model (SAM) [43], introduced by Meta AI, is a large vision transformer [12]-based model trained on a substantial visual corpus. It is a foundation model aiming to address specific downstream image segmentation tasks. One of the challenges in applying deep neural networks to real-world semantic segmentation applications is the need for large amounts of well-annotated training data. SAM effectively addresses this issue by enabling zero-shot generalization to unseen images and objects based on user-provided prompts. The framework of SAM has three main components: an image encoder, a prompt encoder, and a mask decoder as shown in Fig. 2 (a). The image encoder utilizes a vision transformer-based approach for extracting image features. The prompt encoder incorporates user interactions for segmentation tasks. Different types of prompts are supported, including mask, points, box, and text prompts. The mask decoder comprises transformer layers with dynamic mask prediction heads and an Intersection-over-Union (IoU) score regression head. It maps the encoder embedding, prompt embeddings, and an output token to a mask. For user convenience, SAM offers a set of Application Programming Interfaces (APIs) with which segmentation masks can be obtained in just a few lines of code. Different segmentation mode options, e.g., fully automatic, bounding box, and point mode, are supported for different prompts in APIs.
Currently, SAM has made significant strides in diverse fields. In the realm of medical image processing, [44] introduced a straightforward image enhancement method by combining SAM-generated masks with raw images. Additionally, nnSAM [45] integrated the encoder of UNet with the pre-trained SAM encoder, harnessing the feature extraction capabilities of this large foundational vision model. Furthermore, [46, 47] explored SAM’s capacity to generate pseudo labels. Through these distinct technical approaches, these methods have propelled the application and development of SAM across various domains.
@@ II-B Segment Anything Model in Remote Sensing
Figure 3: (a) The schematic diagram of the general semantic segmentation model. (b) The schematic diagram of the proposed method. In our method, the image $X$ is fed into the semantic segmentation model that generates the segmentation output $P$ before the segmentation loss $L_{seg}$ is computed by segmentation output and ground truth $Y$, which is used for back-propagation to update the model. Meanwhile, two additional loss functions $L_{obj}$ and $L_{bdy}$ are computed by SGO $Y_{o}$ and SGB $Y_{b}$ that will also be used for model optimization together with $L_{seg}$.
II-C Object-based methods in Remote Sensing
The semantic segmentation of remote sensing imagery generally revolves two primary approaches: pixel-based and object-based methods [50] that leverage distinct scales for feature learning and final category prediction. There have been many works explored in integrating object-based concepts into various remote sensing tasks [51, 52, 51, 53, 54, 55]. In particular, OCNN [51] proposed the first object-based CNN framework to perform land use classification in the complicated scenarios. SDNF [56] and ESCNet [57] constructed semantic segmentation networks based on superpixels. The former introduced a superpixel-enhanced region module to mitigate noise and strengthen ground object edges, while the latter proposed an adaptive superpixel merging module, optimizing the model towards objects through handling high-dimensional features. OBIC-GCN [58] delved into relationships between objects leveraging graph convolutional networks [59]. However, a notable issue with these methods is that they predominantly focus on extracting object-based features. This approach necessitates the creation of specialized learnable modules to assist in achieving accurate semantic segmentation, thereby elevating the complexity and potential instability of the implementation process [60]. In contrast, the proposed approach integrates seamlessly with existing semantic segmentation models, allowing for straightforward optimization. This method efficiently leverages the raw output of SAM, which contains detailed object information, enhancing overall model performance.
@@ III Methodology
The schematic diagram of the proposed framework is depicted in Fig. 3. Fig. 3 (a) showcases the conventional approach to semantic segmentation, where the input image is fed into the semantic segmentation model to generate segmentation output. This output is then utilized to calculate the segmentation loss, followed by model updates through back-propagation. In contrast, our method, as shown in Fig. 3 (b), incorporates an extra stage that employs the SAM. Specifically, we directly create SGO and SGB using the SAM as provided by Meta AI [25]. These outputs play a crucial role in computing the object consistency loss and boundary preservation loss, respectively, thereby contributing to the model’s training. In this section, we will provide detailed explanations of SAM-based pre-processing, network training, and the associated loss functions.
Figure 4: (a) SGO and its diagram with pixel values, (b) SGB and its diagram with pixel values.
III-A SAM-based Pre-processing
The schematic of the proposed SAM-based pre-processing approach is presented as Fig. 2 (b). SAM provides a grid prompt technique to automatically process images. Given the input remote sensing image denoted by $X\in\mathbb{R}^{H\times W\times 3}$, SAM can generate segmentation masks across the entire image at all plausible locations in the grid prompt setting [44]. In this study, we refer to the segmentation mask as the object, considering each segmentation mask as an individual enclosed region that can be regarded as an object. The generated objects are then stored in a list where we set a threshold of $K$ to limit the maximum number of objects in $X$. Meanwhile, we also establish a threshold $S$ to limit the number of pixels that a single object can contain, effectively filtering out very small segmentation masks. Consequently, an SGO denoted by $Y_{o}\in\mathbb{R}^{H\times W}$ can be obtained, where the value of each pixel falls within the range of $[0,K]$. Pixels not segmented as objects, as well as boundaries, are assigned a value of zero, while the objects in $Y_{o}$ are indexed by an identifier denoted as $i$, where $i\in[1,K]$. The data organization of SGO is presented in Fig. 4 (a). Concurrently, a boundary prior map is derived from SGO. This process involves outlining the exterior boundaries of each object within the list and merging these boundaries to produce a comprehensive boundary prior map, namely SGB, denoted by $Y_{b}\in\mathbb{R}^{H\times W}$. Unless specified otherwise, the identifier for the boundary pixels in $Y_{b}$ is set to $255$, while others are set to 00 in the sequel as shown in Fig. 4 (b). Visual examples of SGO and SGB are illustrated in Fig. 1 (c) and (d).
@@ III-B Network Training
The classical encoder-decoder networks, e.g., UNetformer [21], are widely used in semantic segmentation methods. In this work, we use them as the semantic segmentation model in the proposed framework. Given the input image $X$, the semantic segmentation model produces the predicted segmentation output denoted by $P\in\mathbb{R}^{H\times W\times C}$ where $C$ is the number of categories of ground objects. The learning objective is to minimize the following cross-entropy-based segmentation loss with respect to the parameters of the semantic segmentation model:
@@
where $L_{obj}$ and $L_{bdy}$ stand for the object consistency loss and boundary preservation loss, respectively. Furthermore, $\lambda_{o}$ and $\lambda_{b}$ are two weighting coefficients to balance the three losses.
Figure 5: The flow chart for computing object consistency loss.
III-B1 Object Consistency Loss
The object consistency loss is designed to maintain pixel consistency within the objects in a given input image. Given the input $X$, the output of the semantic segmentation model is denoted as $P$. To compute the object consistency loss, we iterate through all the objects in $Y_{o}$. The data flow is presented in Fig. 5. For each object, we first extract its mask $M^{i}$, i.e. the region in $Y_{o}$ where the pixel value equals $i$. The object feature is then obtained by: